OpenStudy (anonymous):

Five more than a number tripled is half the number How would I make this an equation?

jigglypuff314 (jigglypuff314):

"more than" means adding (+) "a number" can be replaced with x "tripled" means times 3 "is" means equals (=) "half the" means times (1/2) "the number" refers to the x so replace words with it's meanings

OpenStudy (anonymous):

 For "a number" choose a variable to represent it. I chose A. When you do problems like these, just write out an equation while you're saying it. A number tripled would be variable to the third power, five more than is +5 and "is" always means the equal sign. = 1/2(whatever variable you choose to represent "a number" goes here) So it would be A^3 + 5 = 1/2A